OpenStudy (anonymous):

What is the vertex of y=|3x+6|-4 ?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):



OpenStudy (campbell_st):

I would have thought the vertex is where 3x + 6 = 0 or where x = -2 so the vertex is (-2, -4)

OpenStudy (anonymous):

a simple "typo" dude

OpenStudy (phi):

the vertex happens where the stuff inside the | | (absolute value signs) is 0 that means you solve for when 3x+6 = 0 if you know algebra, you would find x=-2 is when 3x+6 is zero once you know x=-2, solve for y using y=|3x+6|-4 of course you already know 3x+6 is zero, so this is y= 0-4 or simply y=-4 the (x,y) pair of the vertex is (-2,-4)